OpenStudy (anonymous):

I really need help! Rewrite the following expressions in x=my + b form. a. 3x + 6y + 9 = 0 b. -2x + 6y - 4 = 0 c. -4x - 2y = 0

OpenStudy (anonymous):

Help!

OpenStudy (amistre64):

x = my form? is that a typo?

OpenStudy (anonymous):

wait! WHAT!???

OpenStudy (amistre64):

a. 3x + 6y + 9 = 0 x = -2y -3 b. -2x + 6y - 4 = 0 x = 3y -2 c. -4x - 2y = 0 x = (-1/2)y

OpenStudy (anonymous):

oh thank! i will discuss and try to understand how to do it :)

OpenStudy (amistre64):

k :)